OpenStudy (anonymous):

1/3 x^=2x+1 solve by an appropriate technique??

OpenStudy (anonymous):

x/3 = 2x + 1 -1 = 2x - x/3 -3 = 6x - x -3 = 5x then x = -3/5 = -0.6

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

What is x^???

OpenStudy (anonymous):

X (squared) to waterineyes

OpenStudy (anonymous):

Then the answer is wrong.

OpenStudy (anonymous):

It will become a quadratic equation and x will have two solutions and not one..

OpenStudy (anonymous):

is it (1/3)*x6 or 1/3x^2

OpenStudy (anonymous):

1/3x^2 sai1234

OpenStudy (anonymous):

\[\frac 13x^2 = 2x + 1\]\[\frac 13 x^2 - 2x - 1 = 0\]At this point, I tend to multiply by 3 to get rid of the fraction.\[3(\frac 13x^2 - 2x - 1 = 0)\]\[x^2 - 6x - 3 = 0\]Can you solve from here?

OpenStudy (anonymous):

Multiply by 3 both the sides: \[x^2 = 6x + 3 \implies x^2 - 6x - 3 = 0\]

OpenStudy (anonymous):

You have to use here quadratic formula: \(D = b^2 - 4ac\) \[x = \frac{-b \pm \sqrt{D}}{2a}\] here a = 1 , b = -6 and c = -3 plug in the values and then find x..

OpenStudy (anonymous):

so where does D come in?

OpenStudy (anonymous):

ahh i see

OpenStudy (anonymous):

D is Discriminant..

OpenStudy (anonymous):

D=b2−4ac x2−6x−3=0 compare with ax^2+bx+c=0 to get a,b,c values

OpenStudy (anonymous):

-6+-square root 48/2 is what i have?